An astronaut, whose height on the earth is exactly 6 ft is lying parallel to the axis of a spacecraft moving at 0.800c relative to the earth. Calculate the height as measured by an on the earth and by an observer on the same same spacecraft.
1
Expert's answer
2020-10-12T05:28:08-0400

For an observer on the Earth height will be calculated by the equation:

"L = L_0\sqrt{1 - \frac{v^2}{c^2}}"

"L = 6\sqrt{1 – 0.64}"

"L = 3.6\;ft"

An observer in the same spacecraft will measure the height to be 6 ft, since there is not relative motion between the observer and the astronaut (v = 0 in the above equation).